### v1

@@ 4 Results
In this Section, we present the reconstruction results of self-calibration algorithm from mock data and DESI LRG sample. Throughout all the analysis, we utilize the $\ell$ modes from $\ell_{\rm{min}}=100$ to $\ell_{\rm{max}}=1000$ and divide into 6 broad bands with edges $\ell=$ 100, 418, 583, 711, 819, 914 and 1000. For the mock datasets, the CCL [^2] https://github.com/LSSTDESC/CCL from [83] is employed to calculate the angular power spectra. Note that the Limber approximation used on $\ell>100$ achieves the modelling accuracy needed for Stage IV surveys [84, 85]. For DESI LRG sample, we use the HEALPix [86] with $N_{\text{side}}=1024$ to construct galaxy overdensity map, and measure the angular power spectrum of galaxies with the function compute_coupled_cell in namaster [87]. When implementing the self-calibration algorithm, we traditionally input $\ell C_{\ell}^{gg,P}$ following the previous work. Nevertheless, for this modified version, multiplying by $\ell$ is no longer necessarily required, as the weighting has already been integrated into the new update rules for NMF. For the mock datasets in Section 4.1, we apply this algorithm to 100 realizations, each with an independent Gaussian random noise generated using equation (2) and identical signal power spectra. The same perturbation procedure is also conducted on the measured angular power spectra from DESI LRG sample in Section 4.2.
@@ 4.1 Results from mock datasets
Despite conducting a range of data tests that varied in photo-$z$ errors, noise level, and redshift bin width, only a subset of these results are showcased in this paper. Initially, we apply the algorithm to the mock data characterized by Low Photo-$z$ Error and both two noise level described in Section 3.2. The redshift range from 0.1 to 1.3 is divided into 20 bins, each with a width of 0.06. The results of biases on scattering matrices ($P^{\rm recon}-P^{\rm true}$) and $\sigma_{\mathrm{NMAD}}$ are shown in Figure 2. We find that accurate reconstructions can be achieved by using the self-calibration algorithm with its newly integrated update rules. The mean absolute bias for all elements in the reconstructed scattering matrix is about 0.011 (0.008) for the case with High Noise (Low Noise). The comparison results demonstrate that the reconstruction accuracy can be obviously enhanced by expanding the area and increasing the number density.
@@
(iii) we extend the region from both sides, ensuring that the average value of the sum of the diagonal and sub-diagonal elements from the newly added two photo-$z$ bins falls within the range $[\mu_{c}-\sigma_{c},\,\mu_{c}+3\sigma_{c}]$.
@@
Considering the self-calibration method is more robust and applicable when applied to thin redshift bins. In Figure 3 we present the results with the redshift bin widths adjusted to 0.04, while keeping all other aspects of the analysis unchanged. The mean absolute bias across all elements of the reconstructed scattering matrix stands at approximately 0.011 (0.006) for High Noise (Low Noise), which is further reduced to around 0.008 (0.004) within the selected region. From the comparison between the upper (High Noise) and lower (Low Noise) panels, we find that the impact of noise is significantly greater here, which could potentially lead to failures in reconstruction within a thinner binned scheme. When implementing the self-calibration algorithm, we can minimize the redshift bin size to the greatest extent possible until the reconstruction result become messy (instead of distributed around the diagonal region) or the algorithm encounters a crash. One point to consider is that the redshift bin width generally needs to exceed the typical photo-$z$ error (such as 0.03 for the cases with Low Photo-$z$ Error), as a distribution that is too diffuse can hinder the accurate reconstruction process. The accuracy can also be self-checked by combing the reconstructed results from different binning schemes to some identical tomographic bins.
Figure 4 demonstrates the reconstructions on the cases with High Photo-$z$ Error. In Appendix Appendix.2 we investigate the ability of the algorithm in the context of a scenario involving a catastrophic overall offset in photo-$z$ error. The still stable performance of results shows the robust capability of the self-calibration algorithm to handle a variety of complex conditions.
@@ 4.2 Application to DESI LRG sample
Reconstructions on the photo-$z$ bins near the redshift edges might be less accurate, but this is unlikely to significantly impact the redshift distribution of the full DESI LRG sample due to the small number of galaxies within those bins. By combining corresponding photo-$z$ bins from current binning scheme, Figure 6 presents the reconstructed results of five samples and the biases in relation to the benchmark references in [70]. The mean absolute bias from the references on the reconstructed results is about 0.008. As previously noted, the $P_{ij}$ scatter carries the same implications as the redshift distribution. To better visualize our results, we show the comparisons in the view of redshift distributions in Figure 7. It is obvious that the self-calibration algorithm can reconstruct accurate redshift distributions using only photometric galaxy clustering, which are comparable to the benchmark references using spectroscopic redshifts from the DESI SV1 data. In order to illustrate the generality and robustness, in Appendix Appendix.3 we also present the reconstructed redshift distributions using another redshift binning scheme with the majority of bin widths set to 0.06.
@@ 5 Discussion and conclusions
[…]
The authors declare that they have no conflict of interest.
Appendix Appendix
Appendix.1 Comparison with previous algorithm
@@
Appendix.2 Mock data with an overall offset redshift error
@@
Appendix.3 DESI LRG sample with another binning scheme
Figure 10: Similar to Figure 7, but for another redshift binning scheme with the majority of bin widths set to 0.06.

### v3

@@ 4 Results
In this section, we present the reconstruction results of self-calibration algorithm from mock data and DESI LRG sample. Throughout all the analysis, we utilize the $\ell$ modes from $\ell_{\rm{min}}=100$ to $\ell_{\rm{max}}=1000$ and divide into 6 broad bands with edges $\ell=$ 100, 418, 583, 711, 819, 914 and 1000. For the mock datasets, the CCL [^3] https://github.com/LSSTDESC/CCL from [83] is employed to calculate the angular power spectra. Note that the Limber approximation used on $\ell>100$ achieves the modelling accuracy needed for Stage IV surveys [84, 85]. For DESI LRG sample, we use the HEALPix [86] with $N_{\text{side}}=1024$ to construct galaxy overdensity map, and measure the angular power spectrum of galaxies with the function compute_coupled_cell in namaster [87]. When implementing the self-calibration algorithm, we traditionally input $\ell C_{\ell}^{gg,P}$ following the previous work. Nevertheless, for this modified version, multiplying by $\ell$ is no longer necessarily required, as the weighting has already been integrated into the new update rules for NMF. For the mock datasets in section 4.1, we apply this algorithm to 100 realizations, each with an independent Gaussian random noise generated using eq. (2.2) and identical signal power spectra. The same perturbation procedure is also conducted on the measured angular power spectra from DESI LRG sample in section 4.2.
@@ 4.1 Results from mock datasets
Table 1: Summary of the mean redshift biases in the photometric redshift bins for different mock datasets. The five equal-width sequential photometric redshift bins have widths set to 0.12, and the total redshift ranges for the cases in figure 2 and figure 3 are [0.40, 1.00] and [0.42, 1.02], respectively.
Case Mean redshift bias $|\Delta z|/(1+z)$
Bin 1 Bin 2 Bin 3 Bin 4 Bin 5
Figure 2 High Noise $0.004$ $0.005$ $0.007$ $0.004$ $0.0007$
Low Noise $0.003$ $0.003$ $0.005$ $0.004$ $0.0005$
Figure 3 High Noise $0.005$ $0.007$ $0.008$ $0.007$ $0.002$
Low Noise $0.004$ $0.004$ $0.004$ $0.003$ $0.001$
Despite conducting a range of data tests that varied in photo-$z$ errors, noise level, and redshift bin width, only a subset of these results are showcased in this paper. Initially, we apply the algorithm to the mock data characterized by Low Photo-$z$ Error and both two noise level described in section 3.2. The redshift range from 0.1 to 1.3 is divided into 20 bins, each with a width of 0.06. The results of biases on scattering matrices ($P^{\rm recon}-P^{\rm true}$) and $\sigma_{\mathrm{NMAD}}$ are shown in figure 2. We find that accurate reconstructions can be achieved by using the self-calibration algorithm with its newly integrated update rules. The mean absolute bias across all elements of the reconstructed scattering matrix is approximately 0.011 (0.008) for the case with High Noise (Low Noise). The comparison results demonstrate that the reconstruction accuracy can be obviously improved by expanding the area and increasing the number density. Consequently, we anticipate that the reconstruction results for stage-IV cosmic shear samples will exhibit enhanced accuracy.
@@
(iii) we extend the region from both sides, ensuring that the average value of the sum of the diagonal and sub-diagonal elements from the newly added two photo-$z$ bins falls within the range $[\mu_{c}-\sigma_{c},\,\mu_{c}+3\sigma_{c}]$. Due to the typically large reconstruction uncertainties in low-density redshift bins, particularly in regions close to the redshift edges, the algorithm has a relatively high probability of finding a local minimum solution, resulting in a rapid decrease in $\mu_{c}$. Thus, we set this stringent lower limit to prevent issues that could arise from any potential local minima.
@@
In order to demonstrate the reconstruction performance more adequately and to compare with the requirements of surveys, we consider five equal-width sequential test photometric redshift bins within $1.00.4\leq z_{p}<$ 1.0. The widths of these test redshift bins are set to 0.12 to match the practical data analysis in surveys, where each bin width typically exceeds 0.1. We note that this redshift range is within the selected region. Then we combine the corresponding photo-$z$ bins from the current thin binning scheme of the reconstructed data, and obtain the redshift distributions of five broad photo-$z$ bins. We present the mean redshift biases for these five photo-$z$ bins in table 1, and it is obvious that the mean redshift of each bin can be accurately reconstructed with the self-calibration algorithm. Although the reconstruction accuracy may fall short of the requirements for the stage-IV surveys (e.g., $0.002(1+z)$ for LSST Year 1 weak lensing analysis), it is comprehensible given the much higher noise levels.
Considering the self-calibration method is more robust and applicable when applied to thin redshift bins. In figure 3 we present the results with the redshift bin widths adjusted to 0.04, while keeping all other aspects of the analysis unchanged. The mean absolute bias across all elements of the reconstructed scattering matrix stands at approximately 0.011 (0.006) for High Noise (Low Noise), which is further reduced to around 0.008 (0.004) within the selected region. Furthermore, the mean redshift biases for five equal-width test photometric redshift bins (within $1.020.42\leq z_{p}<$ 1.02) are also presented in table 1. Here, the total redshift range varies slightly due to the binning scheme adopted. We find that the impact of noise is significantly greater here, which could potentially lead to failures in reconstruction within a thinner binned scheme. When implementing the self-calibration algorithm, we can minimize the redshift bin size to the greatest extent possible until the reconstruction result become messy (instead of distributed around the diagonal region) or the algorithm encounters a crash. The accuracy can also be self-checked by combing the reconstructed results from different binning schemes to some identical tomographic bins.
Figure 4 demonstrates the reconstructions on the cases with High Photo-$z$ Error. In appendix B we investigate the ability of the algorithm in the context of a scenario involving a catastrophic overall offset in photo-$z$ error. The still stable performance of results shows the robust capability of the self-calibration algorithm to handle a variety of complex conditions.
@@ 4.2 Application to DESI LRG sample
Reconstructions on the photo-$z$ bins near the redshift edges might be less accurate, but this is unlikely to significantly impact the redshift distribution of the full DESI LRG sample due to the small number of galaxies within those bins. By combining corresponding photo-$z$ bins from current binning scheme, figure 6 presents the reconstructed results of five samples and the biases in relation to the benchmark references in [70]. The mean absolute bias from the references on the reconstructed results in figure 6 is about 0.008. As previously noted, the $P_{ij}$ scatter carries the same implications as the redshift distribution. To better visualize our results, we show the comparisons in the view of redshift distributions in figure 7. It is obvious that the self-calibration algorithm can reconstruct accurate redshift distributions using only photometric galaxy clustering, which are comparable to the benchmark references using spectroscopic redshifts from the DESI SV1 data. The biases of mean redshifts for full DESI LRG sample and four subsamples between the reconstructions and references are $\sim 0.012(1+z),~{}\sim 0.005(1+z),~{}\sim 0.002(1+z),~{}\sim 0.003(1+z),~{}\sim 0.014(1+z)$, respectively. Note that there are inevitable uncertainties in the references due to the inherent limitations of the spectroscopic redshift data. We anticipate that combining the self-calibration method with other statistics, as detailed in the subsequent section, will significantly enhance accuracy and reduce uncertainty. In order to illustrate the generality and robustness, in appendix C we also present the reconstructed redshift distributions using another redshift binning scheme with the majority of bin widths set to 0.06.
@@ 5 Discussion and conclusions
Appendix A Comparison with previous algorithm
@@
Appendix B Mock data with an overall offset redshift error
@@
Appendix C DESI LRG sample with another binning scheme
Figure 10: Same as figure 7, but for another redshift binning scheme with the majority of bin widths set to 0.06.